OpenStudy (anonymous):

People with type O-negative blood are universal donors. That is, any patient can receive a transfusion of O-negative blood. Only 8.5% of the American population have O-negative blood. If 12 people appear at random to give blood, what is the probability (±0.0001) that at least one of them is a universal donor?

OpenStudy (kropot72):

The probability of one person selected at random of having type O-negative blood is\[\frac{85}{1000}\] The probability that at least one of 12 having type O-negative blood is \[\frac{85\times 12}{1000}=1.02\] The maximum possible value of probability is 1, therefore the answer is P = 1